# Data-Free Continual Learning of Server Models in Model-Heterogeneous Federated learning

### v1

@@ V-A1 Dataset setting
We conduct experiments on two benchmark datasets, Grayscale and RGB, to evaluate various methods. Specifically, Grayscale benchmark is constructed by aggregating the first 10 classes from each of MNIST[24], EMNIST[8], and Fashion-MNIST[54], forming a 30-class grayscale image classification dataset. The RGB benchmark comprises the first 30 classes of CIFAR-100[23], offering a 3-channel image classification dataset.
To emulate a class-incremental continual learning scenario, we divide the label space of each dataset into multiple non-overlapping tasks of equal size, where each task is learned sequentially without revisiting previous classes. Here, we set this to 3 tasks. Furthermore, to reflect practical data heterogeneity among clients within each task, we adopt a non-independent and identically distributed (non-IID) data partitioning scheme. The extent of data heterogeneity is described by the Dirichlet distribution $Dir(\gamma)$[19], where a smaller value of $\gamma$ corresponds to a higher degree of non-IIDness. In our experiments, we evaluate two heterogeneity levels by setting $\gamma=0.5$ and $\gamma=0.3$, respectively.
@@ V-A2 Model setting
We consider a federated learning scenario in which heterogeneous client models collaboratively and continuously train a larger server-side model. Client models are constructed according to their varying resources. In our experiments, we follow the principal works in FCL[64][31][45], employing a total of five clients. Based on their parameter scales, we categorize the client models into three types: Large (L), Medium (M), and Small (S). Client 1, possessing the abundant resources, is assigned an L-type model. Clients 2 and 3, with moderate computational capacity, are equipped with M-type models. Clients 4 and 5, facing the constrained resource budgets, are assigned S-type models.
For Grayscale, a 4-layer CNN is adopted as both the L-type client model and the server model, while two 2-layer CNNs with different convolutional configurations are used as the M-type and S-type models, respectively. For RGB, ResNet-50, ResNet-32, and ResNet-18[16] are assigned as the L-type, M-type, and S-type models, respectively. On the server side, to reflect architectural diversity, we employ a pre-trained Vision Transformer (ViT)[11] and fine-tune only its classification head. The detailed parameter information is presented in Figure 5.
@@ V-A3 Baselines
We first consider baseline methods that assume homogeneous client models. These include the original FedAvg[37], federated variants of classic continual learning approaches such as FedEWC [25] and FLwF-2T[46] inspired by LWF[30], as well as latest data-free FCL methods that rely on synthetic data, such as Target[64] and Lander[45]. For all these homogeneous baselines, we configure the models on all five clients as S-type. Noted that Lander relies on dataset-specific pre-trained label embeddings and is therefore inapplicable to our first dataset.
To evaluate under model heterogeneity, we further include heterogeneous-model approaches. Specifically, we adopt representative federated knowledge distillation KD strategies[5][26] as the foundational communication mechanism for heterogeneous models, and combine them with brief fine-tuning as well as classic continual learning methods, EWC and LwF, to form additional hybrid baselines (KD + FT/EWC/LWF). In all heterogeneous methods, the client and server model architectures are configured identically to those in our proposed framework.
@@ V-A4 Evaluation metrics
Following the previous FCL works[64][31][45], we evaluate performance with two evaluation metrics: accuracy (ACC) and forgetting measure (F)[4]. specifically, we report the server’s accuracy on previously accumulated tasks after each task to evaluate different methods. In addition, we measure forgetting after the final task. The average forgetting at k-th task can be defined as:
@@ V-A5 Implementation
For Grayscale, We configure 5 pototype update epoch and 50 local steps with 128 batch size. We set $\mathcal{D}^{\text{syn}}_{\text{cur}}$, $\mathcal{D}^{\text{syn}}_{\text{pre}}$, $\mathcal{D}^{\text{syn'}}_{\text{cur}}$, $\mathcal{D}^{\text{syn'}}_{\text{pre}}$ to 200, 100, 500 and 500, respectively. We set rounds $Q$, $Q_{l}$, $Q_{k\rightarrow s}$, $Q_{s\rightarrow k}$ and batch size to 5, 1, 2, 1 and 128, respectively. The pre-trained diffusion model comes from the minDiffusion  [^1] https://github.com/byrkbrk/conditional-ddpm trained on Mnist, Emnist and Fashion-Mnist datasets after 100 epochs. For RGB, We configure 10 pototype update epoch and 50 local steps with 12 batch size. We set $\mathcal{D}^{\text{syn}}_{\text{cur}}$, $\mathcal{D}^{\text{syn}}_{\text{pre}}$, $\mathcal{D}^{\text{syn'}}_{\text{cur}}$, $\mathcal{D}^{\text{syn'}}_{\text{pre}}$ to 200, 50, 200 and 200, respectively. We set rounds $Q$, $Q_{l}$, $Q_{k\rightarrow s}$, $Q_{s\rightarrow k}$ and batch size to 40, 5, 1, 1 and 128, respectively. The pre-trained diffusion model comes from the latent diffusion model (LDM)   [^2] https://github.com/CompVis/latent-diffusion[rombach2021highresolution].
@@ V-B Performance Evaluation
[…]
@@
$D_{\text{cur}}^{\text{syn}}$
in ②
$D_{\text{pre}}^{\text{syn}}$
in ②
server
accuracy
client
accuracy
@@ VI Conclusion
In this study, we proposed FedDCL, a novel federated framework that pioneers the continual collaboration of heterogeneous clients in training a server model within a knowledge stream, all in a data-free manner. Through federated prototype extraction, we leveraged pre-trained diffusion models to obtain lightweight class prototypes that capture the knowledge of the current task. These prototypes effectively facilitate data-free knowledge transfer and data-free knowledge retention. Through data augmentation and data replay, we enhance the continuity and stability of local training, and by employing a multi-teacher distillation strategy, we effectively transfer knowledge to the server model, enabling its continual updating. Compared with the state-of-the-arts, our algorithm improved the server accuracy by 9.00%/6.48%/24.37%/18.96% on four settings.

### v2

@@ V-A1 Dataset setting
We conduct experiments on two benchmark datasets, Grayscale and RGB, to evaluate various methods. Specifically, Grayscale benchmark is constructed by aggregating the first 10 classes from each of MNIST[lecun1998gradient], EMNIST[cohen2017emnist], and Fashion-MNIST[xiao2017fashion], forming a 30-class grayscale image classification dataset. The RGB benchmark comprises the first 30 classes of CIFAR-100[krizhevsky2009learning], offering a 3-channel image classification dataset.
To emulate a class-incremental continual learning scenario, we divide the label space of each dataset into multiple non-overlapping tasks of equal size, where each task is learned sequentially without revisiting previous classes. Here, we set this to 3 tasks. Furthermore, to reflect practical data heterogeneity among clients within each task, we adopt a non-independent and identically distributed (non-IID) data partitioning scheme. The extent of data heterogeneity is described by the Dirichlet distribution $Dir(\gamma)$[hsu2019measuring], where a smaller value of $\gamma$ corresponds to a higher degree of non-IIDness. In our experiments, we evaluate two heterogeneity levels by setting $\gamma=0.5$ and $\gamma=0.3$, respectively.
@@ V-A2 Model setting
We consider a federated learning scenario in which heterogeneous client models collaboratively and continuously train a larger server-side model. Client models are constructed according to their varying resources. In our experiments, we follow the principal works in FCL[target][dddr][lander], employing a total of five clients. Based on their parameter scales, we categorize the client models into three types: Large (L), Medium (M), and Small (S). Client 1, possessing the abundant resources, is assigned an L-type model. Clients 2 and 3, with moderate computational capacity, are equipped with M-type models. Clients 4 and 5, facing the constrained resource budgets, are assigned S-type models.
For Grayscale, a 4-layer CNN is adopted as both the L-type client model and the server model, while two 2-layer CNNs with different convolutional configurations are used as the M-type and S-type models, respectively. For RGB, ResNet-50, ResNet-32, and ResNet-18[he2016deep] are assigned as the L-type, M-type, and S-type models, respectively. On the server side, to reflect architectural diversity, we employ a pre-trained Vision Transformer (ViT)[vit] and fine-tune only its classification head. The detailed parameter information is presented in Figure 5.
@@ V-A3 Baselines
We first consider baseline methods that assume homogeneous client models. These include the original FedAvg[Fedavg], federated variants of classic continual learning approaches such as FedEWC [lee2017overcoming] and FLwF-2T[fedLWF2T] inspired by LWF[li2017learning], as well as latest data-free FCL methods that rely on synthetic data, such as Target[target] and Lander[lander]. For all these homogeneous baselines, we configure the models on all five clients as S-type. Noted that Lander relies on dataset-specific pre-trained label embeddings and is therefore inapplicable to our first dataset.
To evaluate under model heterogeneity, we further include heterogeneous-model approaches. Specifically, we adopt representative federated knowledge distillation KD strategies[fedgems][li2019fedmd] as the foundational communication mechanism for heterogeneous models, and combine them with brief fine-tuning as well as classic continual learning methods, EWC and LwF, to form additional hybrid baselines (KD + FT/EWC/LWF). In all heterogeneous methods, the client and server model architectures are configured identically to those in our proposed framework.
@@ V-A4 Evaluation metrics
Following the previous FCL works[target][dddr][lander], we evaluate performance with two evaluation metrics: accuracy (ACC) and forgetting measure (F)[forgettingmeasure]. specifically, we report the server’s accuracy on previously accumulated tasks after each task to evaluate different methods. In addition, we measure forgetting after the final task. The average forgetting at k-th task can be defined as:
@@ V-A5 Implementation
For Grayscale, We configure 5 pototype update epoch and 50 local steps with 128 batch size. We set $\mathcal{D}^{\text{syn}}_{\text{cur}}$, $\mathcal{D}^{\text{syn}}_{\text{pre}}$, $\mathcal{D}^{\text{syn'}}_{\text{cur}}$, $\mathcal{D}^{\text{syn'}}_{\text{pre}}$ to 200, 100, 500 and 500, respectively. We set rounds $Q$, $Q_{l}$, $Q_{k\rightarrow s}$, $Q_{s\rightarrow k}$ and batch size to 5, 1, 2, 1 and 128, respectively. The pre-trained diffusion model comes from the minDiffusion  [^1] https://github.com/byrkbrk/conditional-ddpm trained on Mnist, Emnist and Fashion-Mnist datasets after 100 epochs. For RGB, We configure 10 pototype update epoch and 50 local steps with 12 batch size. We set $\mathcal{D}^{\text{syn}}_{\text{cur}}$, $\mathcal{D}^{\text{syn}}_{\text{pre}}$, $\mathcal{D}^{\text{syn'}}_{\text{cur}}$, $\mathcal{D}^{\text{syn'}}_{\text{pre}}$ to 200, 50, 200 and 200, respectively. We set rounds $Q$, $Q_{l}$, $Q_{k\rightarrow s}$, $Q_{s\rightarrow k}$ and batch size to 40, 5, 1, 1 and 128, respectively. The pre-trained diffusion model comes from the latent diffusion model (LDM)   [^2] https://github.com/CompVis/latent-diffusion[rombach2022highresolutionimagesynthesislatent].
@@ V-B Performance Evaluation
The experimental data shows that our proposed FedDCL outperforms all baselines across all settings in terms of both cumulative accuracy after learning all tasks (-$t_{3}$, higher is better), and forgetting measure (F, lower is better). Specially, on the Grayscale datasets under $Dir(\gamma=0.3)$, our approach attains 80.92% cumulative accuracy, 9.00 percentage points higher than the second-best method Target’s 71.92, while reducing forgetting to 4.37, significantly outperforming all baselines. Similar gains are observed under $Dir(\gamma=0.5)$, where we achieve 82.99% and 4.86, surpassing the runner-up by 6.48 percentage points in accuracy and cutting forgetting by 6.66. On the more challenging RGB datasets, our method again leads by a large margin. We obtain 74.10% and 73.43% cumulative accuracy, improving 23.67% and 16.40% compared to the closest baseline Lander, and we also achieve the lowest forgetting measure 8.40 and 6.85, respectively. This is because our approach incorporates heterogeneous models and leverages the advantages of server-side ViT fine-tuning. This highlights the effectiveness of FedDCL in training the server model continuously on task flow.
@@
$D_{\text{cur}}^{\text{syn}}$ in ② $D_{\text{pre}}^{\text{syn}}$ in ② server accuracy client accuracy
@@ VI Conclusion
In this study, we proposed FedDCL, a novel federated framework that pioneers the continual collaboration of heterogeneous devices in training a cloud server model within a knowledge stream, all in a data-free manner. Through federated prototype extraction, we leveraged pre-trained diffusion models to obtain lightweight class prototypes that capture the knowledge of the current task. These prototypes effectively facilitate data-free knowledge transfer and data-free knowledge retention. Through data augmentation and data replay, we enhance the continuity and stability of local training, and by employing a multi-teacher distillation strategy, we effectively transfer knowledge to the server model, enabling its continual updating. Compared with the state-of-the-arts, our algorithm improved the server accuracy by 9.00%/6.48%/23.67%/16.40% on four settings.
[Uncaptioned image] Xiao Zhang is now an associate professor in the School of Computer Science and Technology, Shandong University. His research interests include data mining, distributed learning and federated learning. He has published more than 30 papers in the prestigious refereed journals and conference proceedings, such as IEEE Transactions on Knowledge and Data Engineering, IEEE Transactions on Mobile Computing, NeurIPS, SIGKDD, SIGIR, UBICOMP, INFOCOM, ACM MULTIMEDIA, IJCAI, and AAAI.
[Uncaptioned image] Zengzhe Chen is an M.D. student in the School of Computer Science and Technology at Shandong University. He received his B.S. degree from Shandong University. His current research interests include distributed learning.
[Uncaptioned image] Yuan Yuan received the BSc degrees from the School of Mathematical Sciences, Shanxi University in 2016, and the Ph.D. degree from the School of Computer Science and Technology, Shandong University, Qingdao, China, in 2021. She is currently a postdoctoral fellow at the Shandong University-Nanyang Technological University International Joint Research Institute on Artificial Intelligence, Shandong University. Her research interests include distributed computing and distributed machine learning.
[Uncaptioned image] Yifei Zou received the B.E. degreefrom Computer School, Wuhan University, Wuhan,China, in 2016, and the Ph.D. degree from the Department of Computer Science, The University of Hong Kong, Hong Kong, in 2020. He is currently an Assistant Professor with the School of Computer Science and Technology, Shandong University, Qingdao, China. His research interests include wireless networks, ad hoc networks, and distributed computing.
[Uncaptioned image] Fuzhen Zhuang received the Ph.D. degrees in the Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, in 2011. He is currently a full Professor with the Institute of Artificial Intelligence, Beihang University. He has published more than 150 papers in some prestigious refereed journals and conference proceed-ings such as Nature Comunications, the IEEE Transactions on Knowledge and Data Engineering, the IEEE Transactions on Cybernetics, the IEEE Transactions on Neural Networks and Learning Systems, the ACM Transactions on Knowledge Discovery from Data, the ACM Transactions on Intelligent Systems and Technology, Information Sciences, Neural Networks, SIGKDD, IJCAI, AAAI, TheWebConf, ACL, SIGIR, ICDE, ACM CIKM, ACM WSDM, SIAM SDM, and IEEE ICDM. His research interests include transfer learning, machine learning, data mining, multitask learning, knowledge graph and recommendation systems. He is a Senior Member of CCF. He was the recipient of the Distinguished Dissertation Award of CAAI in 2013.
[Uncaptioned image] Wenyu Jiao is a Ph.D. student in Finance at the Desautels Faculty of Management, McGill University. She received her B.A. degree in Mathematics and Economics from New York University and her M.S. degree in Financial Mathematics from Johns Hopkins University. Her current research interests include asset pricing, digital assets, data automation, and the application of large language models for data analytics and decision making.
[Uncaptioned image] Yuke Wang is a Ph.D. student in Information System at the Desautels Faculty of Management, McGill University. He received his B.A. degree from Xi’an Jiaotong University and his Master of information studies degree from McGill University. His current research interests include the economic and societal impacts of large language models, IT innovation, and AI transformation.
[Uncaptioned image] Dongxiao Yu received the B.S. degree in 2006 from the School of Mathematics, Shandong University and the Ph.D degree in 2014 from the Department of Computer Science, The University of Hong Kong. He became an associate professor in the School of Computer Science and Technology, Huazhong University of Science and Technology, in 2016. He is currently a professor in the School of Computer Science and Technology, Shandong University. His research interests include edge intelligence, distributed computing and data mining.